OpenStudy (anonymous):

You are given an annual interest rate and the compounding period. Find the interest rate per compounding period. 13.2%; monthly

OpenStudy (anonymous):

is it 1.6%?

OpenStudy (anonymous):

Alright well first you would use the annual future value formula to find the yearly growth and assume a principle value of 100 for ease of calculation: FV=PV(1+rate)^time=(100)(1+.132)^(1)=113.20 Now using the monthly future value formula you can solve for the monthly rate: FV=PV(1+(rate/compound periods))^(time*compounding periods) And since there are 12 months in a year: 113.20=(100)(1+(rate/12))^(1*12) 1.132=(1+(rate/12))^12 1.010386=1+(rate/12) 0.010386=rate/12 rate=0.12463=12.463% So the monthly compounding rate is 12.463%.

OpenStudy (anonymous):

does all that make sense?

OpenStudy (anonymous):

yes thanks!

OpenStudy (anonymous):

no prob